Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
January 25, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Michael Howell, Michael Jay, Bobbie Meng, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Evan Wack, Stephen Williams
Announcements
  1. Bobbie has a 3-month old cat that needs a home.
  2. Winston started a new committee, BIT (Bakers of IT). The mission is to bake cakes once a month for birthdays. We are looking for volunteers to join.
  3. Grover documentation - need to discuss the procedures for checking in machines and recording items in Grover. The history is incorrect. There needs to be consistency.
  4. Will's puppy is now 19.2 lbs. The older dog is only 13 lbs.
  5. Move to Eastside - nothing has changed. Still on for June. Tape vault inventory is still unresolved. Regarding the issue of parking decals that need to renewed in May: Evan, Josh, Laurie, Trish and Michael Howell will be staying on campus - everyone else will be out at Eastside. However, there is no clarity as to what, if any, decal will be required at the Eastside. Meetings will be held via videoconferencing. Campus is going to pick up Office Communicator, so that should make it easier. All checkout items will remain on campus.
  6. Acquisitions probable move to 13th Street locale - the current phone system in place in that building would remain. The servers would be left behind. Both would be our responsibility.
  7. The College of Education's IT Dept asked for help to manage their laptop checkout in ALEPH. Jason met with them on Friday.
Ongoing IT Updates

Orders
  1. Will is working on a difficult order for LCD monitors for MSL.
Hardware
  1. The first meeting using the video conferencing equipment is going on right now. The system crashed during set-up and had to be reset.
  2. High-end computers in West have been replaced. The old ones have been moved to the vault. Grover is being updated accordingly.
  3. George Hack is coming by the IT Dept today at 11:30 to pick up twelve 17-inch monitors.
  4. Evan had a conference call with Altiris about software virtualization and streaming. Need to also talk with Rice University about their experience in using Altiris.
  5. WIN7 image - VIVO laptops have had a few problems with the touchpad. Desktops - Stephen had one random restart. Could be user error or an update.
  6. Rachel is revisiting focus sound for West 234.
  7. Laptops - all circulation laptops in West and MSL have been reimaged and are working great. We have the Education laptops and they want to update JAVA, but it will break E-Learning. We told Education they need to make a decision. Altiris virtual app should fix these problems.
  8. User Safe Taskforce - details presented to IT during our meeting last week. Bonnie (HR) cancelled a meeting with Michael Howell. He will reschedule.
  9. Public printing - at a standstill. Currently, there is no maintenance contract. UF Purchasing left it out. We met with UF Purchasing and Canon last week. We have another meeting scheduled for this Thursday to put the issues on the table and try to get a solution from Canon.
Software
  1. Electronic Reserves Grover mods - Chris Poehlmann is aware that the ball is in his court.
  2. Room reservation app - Judy has been asking for it, but we are currently waiting on Stacey Ewing. Chris can't proceed until he gets a response. All of this is detailed within the ticket, so all parties should be aware of the status.
  3. I-Phone app - Chris is working on it for Library Hours. Shruti is working Computer Availability. They are both just basically Web pages especially designed for these apps.
  4. Macros - some completed, new ones initiated. There is a new one for retroactive dissertations involving Cataloging and Preservation.
  5. ARL-HR - awaiting mods from HR.
  6. VIVO - lots of hardware received. New version (.9) should be out this Friday from Cornell, Florida and Indiana. Stephen met with Warren Curry about Harvester. Stephen should not be traveling anymore for awhile.
  7. ITS HR request - HR portion is finished. HSS portion is almost done, with the exception of two mods. Shruti is working on it.
Network/Server Admin
  1. Virtualization - waiting for CNS to provision new storage.
  2. Requesting a new SQL server to replace SQL2K3.
  3. Getting access to file server storage. Will require moving Web files and paths. Web server side - not much change, but back end files will change location. Will affect Dreamweaver, Contribute, etc.
  4. As soon as the storage is online, we can start migrating our current VMs: WSUS, EPO and printers on the same box. UFDC won't be done until towards the end. VIVO IP will change, but the server name will stay the same. Stephen mentions the VIVO grant included money to expand the server if necessary. Per Logan, RAM and storage will increase cost.
ILS
  1. Update to the Production Client scheduled for this Sunday.
DLC General Issues
  1. FTP & Abbey are waiting for CNS storage to come online.
  2. JPG2K - quote for Aware software approved.
  3. Streaming video - no action.